40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 30 21:29:36 2018, awade, Craig, DailyProgress, PMC, Pre mode cleaner installed in North path 
    Reply  Thu Feb 1 15:25:58 2018, awade, Craig, DailyProgress, PMC, Update working schematic of PMC servo card (older card) D980352-A_AWade20180201Mods.pdf
       Reply  Sat Feb 3 18:08:26 2018, awade, DailyProgress, PMC, A few more modifications to the PMC north board D980352-A_AWade20180202Mods.pdf
          Reply  Fri Jul 6 23:43:54 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board LowPrefFilter.filPMCLowPrefFilterOpt.ipynbU2Buffer.filU2Stability.ipynbD980352-A_AWade20180201Mods_--AGupta20180707Mods.pdf
             Reply  Mon Jul 9 16:15:06 2018, rana, DailyProgress, PMC, New modifications to the PMC north board 
                Reply  Thu Jul 12 15:18:26 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board D980352-AGupta20180712Mods.pdf
                   Reply  Sat Jul 21 14:53:16 2018, awade, DailyProgress, PMC, New modifications to the PMC north board 
                      Reply  Wed Aug 29 12:01:54 2018, anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems D980352-AGupta20180712Mods.pdfPMC_Servo_LISO_Analysis_Vs_Measured.pdf
                         Reply  Thu Aug 30 12:01:37 2018, rana, Summary, PMC, Summarization of efforts to fix PMC Servo problems 
                         Reply  Mon Sep 24 12:53:26 2018, awade and anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems: external mixer and LP filter 
Message ID: 2209     Entry time: Fri Jul 6 23:43:54 2018     In reply to: 2070     Reply to this: 2210
Author: anchal 
Type: DailyProgress 
Category: PMC 
Subject: New modifications to the PMC north board 

Using buzzer, we found that the north path PMC responds strongly to buzzing at 484Hz. This was coincidentally the zero of one of the filters inside PMC feedback circuit. We tried to suppress the output with a 10Hz pole passive LPF at the output but it didn't work. On opening the box, we found that the actual circuit has different values and different circuit topology than assumed. In particular, the resistor at the output was found to be 30.9k instead of 1k. This I removed and shorted as we would use a more robust external LPF which does not depend on the capacitance of the cable. Secondly, the circuit topology of the filter which has a pole at 2Hz and zero at 482Hz was different than our schematic. It was a poor design with weird values. I implemented liso fitting in python wrapper and used it to optimize the circuit. Input file and jupyter notebook of optimization is attached. Also, last time Andrew changed R9 to 100 Ohms from 10 Ohms and was skeptic about the opamps stability with it. I ran a stability diagnostic for this part of the circuit as well. Input file U2Buffer.fil and jupyter notebook are attached.

Quote:

A few more changes to the PMC north board. Hopefully this is the last.

Changing series resistance at the output

There was some limit to the gain before the PMC loop became unstable.  I realized that I hadn't implemented the 10 Hz pole at the output, that could be the reason.  The resistor in series with the output was only 1 kΩ.  The Noliac NAC2124 PZT on the PMC cavity has a specified capacitance of 510 nF and about 5 meters of BNC cabling should have a capacitance of 0.5 nF (for 101pF/m).  This should give a pole of about 312 Hz, we want to bring this down.  I replaced the 1 kΩ series resistance at the HV output with a 30.9 kΩ to bring to pole down to 10 Hz.  

Schematic is updated below.

Removing short from PD RF input

For whatever reason the AC coupling resistor had been deliberately shorted on the PD RF input to the board (C3).  I removed the short.  It is AC coupled again.

 

Altering the Acromag excitation voltage and divider resistors

I realized running the Acromags off 6 V is a little awkward.  Most power supplies come in other standard values like 5 V, 7.5 V, 9V etc.  I've raised the excitation of the binary channels to 9 V and made some changes to the divider resistors. Resistors on SW1 and SW2 binary engages were changed to equal values of 1790 Ω for R1 and R2 (see option D, attachment, PSL:2058). On the blanking pin used to activate the AD602, equal values for R1 and R2 of 154 Ω were used.  These resistor choices give a high voltage of 4.5 V and a low of 0.66 V.  There is also a margin of error if the power supply used goes a little over voltage.

I've ordered standard barrel 2.1x5.5 mm panel Jacks so that the unit can be powered from a standard 9V plug pack.

 

Attachment 1: LowPrefFilter.fil  99 Bytes  Uploaded Sat Jul 7 00:54:24 2018  | Hide | Hide all
r R11 7.96k IN n1
r R12 79.6k n1 OUT
c C11 0.1u n1 n4
r R13 3301 n4 OUT
op LT1028 LT1028 GND n1 OUT
Attachment 2: PMCLowPrefFilterOpt.ipynb  452 kB  Uploaded Sat Jul 7 00:54:36 2018  | Hide | Hide all
{
 "cells": [
  {
   "cell_type": "code",
   "execution_count": 1,
   "metadata": {
    "collapsed": true
   },
   "outputs": [],
   "source": [
... 554 more lines ...
Attachment 3: U2Buffer.fil  123 Bytes  Uploaded Sat Jul 7 00:54:49 2018  | Hide | Hide all
r R10 49.9 IN GND
r R4 220k GND n3
r R5 220k n3 n2
c C9 2u n3 gnd
r R9 1000 n2 gnd
r R2 9.09k n2 OUT
op U2 LT1028 IN n2 OUT
Attachment 4: U2Stability.ipynb  311 kB  Uploaded Sat Jul 7 00:55:00 2018  | Hide | Hide all
{
 "cells": [
  {
   "cell_type": "code",
   "execution_count": 1,
   "metadata": {
    "collapsed": true
   },
   "outputs": [],
   "source": [
... 411 more lines ...
Attachment 5: D980352-A_AWade20180201Mods_--AGupta20180707Mods.pdf  170 kB  Uploaded Sat Jul 7 00:55:27 2018  | Hide | Hide all
D980352-A_AWade20180201Mods_--AGupta20180707Mods.pdf
ELOG V3.1.3-